OpenStudy (anonymous):

Find the cosine of both angle A and angle B. A. cos A = 5/13; cos B = 12/13 B. cos A = 12/13; cos B = 5/13 C. cos A = 5/12; cos B = 12/5 D. cos A = 12/5; cos B = 5/12

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

cosine = adjacent/hypotenuse cos A = 24/26 or 12/13 cos B = 10/26 or 5/13

OpenStudy (anonymous):

cos is adj/hyp